Jamal has waffles for breakfast every Sunday morning. He tops them with blueberry syrup or maple syrup, but sometimes he uses both or neither of the syrups. 

On one Sunday:

Let X be the event ‘Jamal uses blueberry syrup’

Let Y be the event ‘Jamal uses maple syrup’

It is given that

P(X) = 13/30

P(X | Y) = 1/4

P(X | Y ') = 9/14

Find the probability that Jamal uses both blueberry and maple syrup


1
Expert's answer
2021-01-20T01:58:22-0500

Given that "P(X)=\frac{13}{30}" ,"P(X/Y)=\frac{1}{4},P(X/Y')=\frac{9}{14}"

We have to find "P(X\cap Y)."

As "P(X/Y)=\frac{1}{4}"

"\implies \frac{P(X\cap Y)}{P(Y)}=\frac {1}{4}"

"\implies P(X \cap Y)=\frac {1}{4}.P(Y)" .......(1)

Again, "P(X/Y')=\frac {9}{14}"

"\implies \frac { P(X\cap Y')}{P(Y')}=\frac{9}{14}"

"\implies \frac {P(X\cup Y)-P(Y)}{1-P(Y)}=\frac {9}{14}" [ as "P(X\cup Y)=P(X\cap Y')+P(Y)," where "Y" and "X\cap Y'" are disjoint set ]

"\implies 14.P(X\cup Y)=5.P(Y)+9"

"\implies 14.[P(X)+P(Y)-P(X\cap Y)]=5.P(Y)+9"

"\implies 14.[P(X)+P(Y)-\frac{1}{4}P( Y)]=5.P(Y)+9"

"\implies \frac{11}{2}.P(Y)=9-14.P(X)"

"\implies \frac{11}{2}.P(Y)=9-14.(\frac{13}{30})"

"\implies \frac{11}{2}.P(Y)=9-\frac{91}{15}"

"\implies P(Y)=\frac{8}{15}"

Therefore from (1), we have "P(X \cap Y)=\frac {1}{4}.\frac{8}{15}=\frac{2}{15}"

"\therefore" The required probability that Jamal uses both blueberry and maple syrup is "=\frac{2}{15}"